40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8593     Entry time: Thu May 16 23:48:39 2013     In reply to: 8589     Reply to this: 8600   8602
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Kiwamu's sensing matrix measurement script revived 

Koji locked the PRMI for me, and I took some data.  I haven't finished figuring out what to do with it / writing a processing script.

Here is the data, in a python dictionary (not for you to read, but so that it's here and you can use it later if you want).

{'AS55_Q': [['ErrorBarData0', '-1.60826e-05', '0.000154774'], ['ErrorBarData1', '-1.61949e-05', '-9.69142e-05'], ['ITMs', '-0.134432', '0.00240338'], ['PRM', '0.0525864', '0.145516']], 'REFL55_Q': [['ErrorBarData0', '-0.00088166', '-0.00294315'], ['ErrorBarData1', '0.00298076', '-0.000466507'], ['ITMs', '-0.573825', '-0.0865747'], ['PRM', '1.94537', '0.534968']], 'REFL33_Q': [['ErrorBarData0', '0.000868208', '0.000785702'], ['ErrorBarData1', '-0.00136268', '-0.000288528'], ['ITMs', '-0.0653009', '-0.0112035'], ['PRM', '0.875275', '0.419765']], 'REFL11_I': [['ErrorBarData0', '-0.147347', '0.136075'], ['ErrorBarData1', '0.351823', '0.160556'], ['ITMs', '-12.0739', '-80.1513'], ['PRM', '6991.11', '7073.74']], 'REFL33_I': [['ErrorBarData0', '-0.00100624', '0.00134366'], ['ErrorBarData1', '0.00373581', '0.000783243'], ['ITMs', '-0.399404', '-0.774793'], ['PRM', '67.4138', '68.8886']], 'REFL11_Q': [['ErrorBarData0', '-0.0173368', '0.0141987'], ['ErrorBarData1', '0.100048', '0.0882165'], ['ITMs', '6.46585', '-26.2841'], ['PRM', '1653.42', '1663.96']], 'AS55_I': [['ErrorBarData0', '-1.87626e-05', '2.24596e-05'], ['ErrorBarData1', '-5.46466e-05', '-2.96552e-07'], ['ITMs', '-0.00531763', '0.00130579'], ['PRM', '-0.100501', '-0.0706334']], 'REFL55_I': [['ErrorBarData0', '-0.000774208', '-5.32631e-05'], ['ErrorBarData1', '0.00347621', '0.0025103'], ['ITMs', '-0.115633', '-0.83847'], ['PRM', '72.8058', '74.2347']]}

The structure is that each sensor has some "error bar" measurements, when there was no drive to any optics (I, then Q of the lockin), and then response to different optics' drives (waiting 20sec after turning on the oscillator before making a measurement, since the lockin has 0.1Hz lowpasses.  ).

The amplitude that Kiwamu had of 4000 cts in the LSC lockin was fine for MICH, but made PRCL unlock, so this data was taken with an amplitude of 1000 counts, at a frequency 283.1030 Hz. 

Since this is only barely above the UGF for both MICH and PRCL loops, I also have OLTF information at 283Hz from DTT:  PRCL mag = -1.05264 dB, phase = 24.6933 deg, MICH mag = -8.50951 dB, phase = 31.3948 deg.

I have started writing a script SensMatAnalysis.py in the scripts/LSC directory to do the analysis, but after having talked to Koji, I need to do more thinking to make sure I know what I'm doing.  Stay tuned for actual analysis later.

ELOG V3.1.3-